# SceneTracker: Long-term Scene Flow Estimation Network

### v1

@@ 4.1.1 Datasets
Unlike optical and scene flow tasks, LSFE lacks a dedicated dataset. For example, scene flow datasets such as FlyingThings3D [12] and KITTI2015 [13] only provide dense 3D point correspondences between adjacent frames. Although these datasets also provide 3D video data containing long-term information, the reference points change continuously as the video progresses.
The only dataset potentially supporting LSFE tasks is PointOdyssey [25], initially proposed for the TAP task. PointOdyssey is a large-scale synthetic dataset and data generation framework for the training and evaluating TAP methods. It aims to push the boundaries of what is possible by focusing on lengthy videos featuring realistic movements. PointOdyssey animates flexible characters using real-world motion capture data, constructs 3D environments corresponding to the motion capture settings, and generates camera perspectives using trajectories obtained through structure-from-motion applied to real videos to achieve this naturalism. It introduces variation by randomizing character attributes, motion characteristics, materials, lighting, 3D elements, and environmental effects.
Fortunately, it can selectively output the depth maps corresponding to RGB images and trajectories in the depth dimension, providing opportunities for training and evaluating LSFE methods. We construct an augmented dataset called LSFOdyssey based on the current PointOdyssey [25] dataset to tailor to LSFE tasks and reduce dataset reading time to expedite training. The LSFOdyssey dataset comprises 127,437 training data samples acquired through data augmentation techniques such as random flipping, spatial transformations, and color transformations. Each sample includes a 24-frame 3D video along with 3D trajectories of 256 query points. Furthermore, we ensure that all query points are non-occluded in at least the first four frames and a total of 12 frames in the video. Besides, we evenly partitioned 90 testing data samples without applying data augmentation techniques. Each sample consists of 40 frames and includes 256 query points. The entire LSFOdyssey dataset is approximately 240GB in size.
@@ 4.1.2 Evaluation Metrics
In evaluating the 3D performance of trajectories, we employ the following configuration. We utilize $\delta_{3D}^{x}$s and their average value $\delta_{3D}^{avg}$ to measure the percentage of trajectory points within a threshold distance $x\in\{0.05,0.10,0.20,0.40\}m$ to ground truth. We use median 3D trajectory error $\text{MAE}_{3D}$ and 3D end-point error $\text{EPE}_{3D}$ to measure the distance between the estimated trajectories and ground truth trajectories. Similarly, we utilize metric $\text{Survival}_{3D}^{0.20}$, with a failure threshold set at $0.20m$.
Figure 2: Visualizations of our SceneTracker’s estimation results on the LSFOdyssey test dataset.
@@ 4.1.3 Implementation Details
We use PyTorch to implement our method. We use the AdamW [10] as an optimizer and the One-cycle [16] as the learning rate adjustment strategy. We set $S=16$, $s=8$, $l=4$, $r=3$, $r_{h}=3$, and $r_{w}=4$. We evaluate our method after 4 FIM iterations, i.e., $n=4$, and we set $M=6$ for the Transformer Cell network. We use the NVIDIA 3090 GPU for training and inference.
@@ 4.1.4 Training Processes
Our training includes a training process named Odyssey. We train the model on the LSFOdyssey dataset from scratch in the Odyssey training process. Each data sample learns 24 time frames and 256 query points. The input resolution is $384\times 512$. The batch size is 8. The learning rate is $2\times 10^{-4}$, and the number of training steps is 200K.
@@ 4.2 Visualizations of Results
Following the Odyssey training process, we visualize the estimation results of SceneTracker on the LSFOdyssey test dataset, as shown in Fig. 2. We present the 14 frames of point clouds from the 40-frame test data sample at equal intervals and the RGB images of the first and last frames. The predicted 3D trajectories are depicted in blue on their corresponding point clouds. From Fig. 2, we can see that our method can consistently deliver smooth, continuous, and precise estimation results in the face of the complex motions of the camera and dynamic objects in the scene.
4.3 Comparisons with SF baseline
Table 1: Quantitative results on the 2D metrics within the LSFOdyssey test dataset. All data come from the Odyssey training process.
@@ 4.3 Comparisons with SF baseline
SceneTracker (Ours) 85.33 96.80 2.02
We follow RAFT [18] to design a straightforward scene flow baseline named SF baseline. Since original scene flow methods can only address trajectory prediction problems between two frames, we adopt a chaining approach to handle long videos. Precisely, after calculating the scene flow between the first and second frames, we determine the $uv$ coordinates corresponding to the predicted position of the second frame and index the depth of the second frame with these coordinates to obtain the query point coordinates from the second frame to the third frame, and so on for subsequent frames.
We also train the SF baseline on the LSFOdyssey dataset from scratch to ensure equity. We use the same training settings as the SceneTracker’s Odyssey training process. After training, we compare our SceneTracker with the SF baseline on the LSFOdyssey test dataset. The results are shown in Fig. 3, Table 1 and Table 2.
We can see from Table 1 and Table 2 that our method significantly outperforms the SF baseline across all dataset metrics, both in 2D and 3D domains. Especially on $\text{MAE}_{2D}$ and $\text{MAE}_{3D}$, SceneTracker achieves error reductions of 58.8% and 39.0%, respectively. In contrast to the scene flow approach, our method employs identical query point features as initial templates tracked for each window, mitigating query point drift. This strategy effectively addresses occlusion and out-of-boundary effects, further substantiated in Fig. 3’s solid-box marked regions. The results fully demonstrate our method’s superior performance. The experimental findings substantiate the efficacy of our approach in tackling the LSF challenge.
4.4 Comparisons with TAP baseline
@@
Table 2: Quantitative results on the 3D metrics within the LSFOdyssey test dataset. All data come from the Odyssey training process.
Method $\delta_{3D}^{0.05}\uparrow$ $\delta_{3D}^{0.10}\uparrow$ $\delta_{3D}^{0.20}\uparrow$ $\delta_{3D}^{0.40}\uparrow$ $\delta_{3D}^{avg}\uparrow$ $\text{Survival}_{3D}^{0.20}\uparrow$ $\text{MAE}_{3D}\downarrow$ $\text{EPE}_{3D}\downarrow$
TAP baseline 42.43 57.45 67.52 74.20 60.40 49.37 0.389 0.575
SF baseline 53.22 71.78 84.40 92.67 75.52 83.38 0.123 0.132
SceneTracker (Ours) 64.67 79.88 90.85 96.67 83.02 89.25 0.075 0.081
We observe that our estimated 3D trajectories can be projected onto the image plane to form a continuous 2D trajectory, which serves as the output of the TAP task. For this purpose, we construct a TAP baseline based on SceneTracker. Specifically, the TAP baseline first outputs the image plane $uv$ trajectory, then utilizes the $uv$ trajectory to index into the depth map of each frame, and finally transforms the indexed $uvd$ trajectory into the camera coordinate system, as shown in Eq. 4.
@@
We utilize SceneTracker trained through the Odyssey training process to provide 2D trajectories for the TAP baseline. After training, we compare our SceneTracker with the TAP baseline on the LSFOdyssey test dataset. The results are shown in Fig. 3 and Table 2.
@@
We can see from Table 2 that our approach significantly outperforms the TAP baseline on all 3D dataset metrics. Especially on $\text{MAE}_{3D}$ and $\text{EPE}_{3D}$, SceneTracker achieves error reductions of 80.7% and 85.9%, respectively. At the same time, on $\text{Survival}_{3D}^{0.20}$, SceneTracker achieves an 80.8% increase in accuracy. The significant performance gap stems mainly from the TAP baseline lacking robustness against depth noise interference and the inability to handle occlusion relationships in 3D space. It is worth noting that LSFOdyssey simulates fog and smoke, setting their depths as outliers directly, resulting in a depth map with a substantial amount of noise, posing significant challenges. The TAP baseline, due to its direct indexing from the depth map, is unable to detect and correct erroneous depth data, which is also substantiated in Fig. 3.
@@
The results fully demonstrate that the proposed method excels in spatial relationship handling and noise resistance, highlighting the unique challenges the LSF task poses.
@@ 4.5 Ablation Studies
Table 3: The model is trained with stride 8 and sliding window size $S=16$. The same inference sliding window size gives the best results.
Window Size $\delta_{2D}^{avg}\uparrow$ $\text{Survival}_{2D}^{16}\uparrow$ $\text{MAE}_{2D}\downarrow$ $\delta_{3D}^{avg}\uparrow$ $\text{Survival}_{3D}^{0.20}\uparrow$ $\text{MAE}_{3D}\downarrow$
8 83.17 96.29 2.89 78.35 86.51 0.102
12 84.68 96.51 2.22 81.47 88.08 0.083
16 85.33 96.80 2.02 83.02 89.25 0.075
20 84.48 96.50 2.11 53.21 58.03 0.197
24 83.93 96.31 2.26 44.75 47.80 0.232
We conduct a series of ablation experiments on our method from 2 aspects, which include the inference sliding window size, and the number of FIM’s updates, to examine their contributions. We use the model after Odyssey in all experiments and evaluate the LSFOdyssey test dataset.
4.5.1 Inference sliding window size
In Table 3, we inspect how the sliding window size of a model trained with sliding window $S=16$ affects model performance. We find that an inference window size $S=16$ achieves optimal performance across all dataset metrics, particularly all 3D metrics.
4.5.2 Number of FIM’s updates
Figure 4: Dataset evaluation accuracy versus the number of FIM’s updates $N$ at inference time.
We train the model with $N=4$ FIM’s updates and evaluate with $N=4$. The same inference update number gives the best results as evaluated in Fig. 4. The “Accuracy” represents the metric $\delta_{3D}^{avg}$.
@@ 5 Conclusion
In summary, we propose a new task, i.e., long-term scene flow estimation, to better capture the fine-grained and long-term 3D motion. To effectively tackle this issue, we introduce a novel approach named SceneTracker. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs.

### v2

@@ 5.1.1 Datasets
LSFOdyssey Unlike optical and scene flow tasks, LSFE lacks a dedicated dataset. We reorganize the PointOdyssey [20] dataset initially proposed for the TAP task. PointOdyssey is a large-scale synthetic dataset and data generation framework for the training and evaluation of TAP methods. Fortunately, it can selectively output the depth maps corresponding to RGB images and trajectories in the depth dimension, providing opportunities for training and evaluating LSFE methods. We construct an augmented dataset called LSFOdyssey based on the current PointOdyssey [20] dataset to tailor to the LSFE task and reduce dataset reading time to expedite training. The LSFOdyssey dataset comprises 127,437 training data samples acquired through data augmentation techniques such as random flipping, spatial transformations, and color transformations. Each sample includes a 24-frame 3D video along with 3D trajectories of 256 query points. Furthermore, we ensure that all query points are non-occluded in at least the first four frames and a total of 12 frames in the video. Besides, we evenly partition 90 testing data samples without applying data augmentation techniques. Each sample consists of 40 frames and includes 256 query points. The entire LSFOdyssey dataset is approximately 240GB in size.
LSFDriving Following the Section 4, we construct the first real-world evaluation dataset LSFDriving for the LSFE field. We obtained RGB videos and annotations for backgrounds and vehicles from the Waymo [26] dataset. Due to the absence of binocular cameras in Waymo, we apply depth completion methods to generate dense depth sequences. Specifically, we interpolate the nearest neighbors. It is worth noting that we utilize completed depth maps to establish occlusion relationships, thereby providing valid masks for evaluation. Meanwhile, we acquired RGB videos for pedestrians from the KITTI [29] raw data collection. We utilize LEAStereo [28] to provide dense depth sequences and the proposed semi-automated annotation pipeline to provide annotations. We do not offer annotations for left-view occluded and binocular non-co-visible regions. Ultimately, each category comprises 60 testing data samples consisting of 40 frames. There are 5 query points for each background and vehicle sample. We annotate 1-5 challenging non-rigid motion points in every pedestrian sample, such as joint and garment points.
@@ 5.1.2 Evaluation Metrics
When evaluating the 3D performance of trajectories, we employ the following configuration. We utilize $\delta_{3D}^{x}$s and their average value $\delta_{3D}^{avg}$ to measure the percentage of trajectory points within a threshold distance $x\in\{0.10,0.20,0.40,0.80\}m$ to ground truth. We use median 3D trajectory error $\text{MAE}_{3D}$ and 3D end-point error $\text{EPE}_{3D}$ to measure the distance between the estimated trajectories and ground truth trajectories. Similarly, we utilize metric $\text{Survival}_{3D}^{0.50}$, with a failure threshold set at $0.50m$.
@@ 5.1.3 Implementation Details
We use PyTorch to implement our method. We use the AdamW [30] as an optimizer and the One-cycle [31] as the learning rate adjustment strategy. We set $S=16$, $s=8$, $l=4$, and $r=3$. We evaluate our method after 4 FIM iterations, i.e., , $n=4$, and we set $M=6$ for the Transformer Cell network. We use the NVIDIA 3090 GPU for training and inference.
@@ 5.1.4 Training Processes
Our training includes a training process named Odyssey. In the Odyssey training process, we train the model on the LSFOdyssey dataset from scratch. Each data sample learns 24 time frames and 256 query points. The input resolution is $384\times 512$. The batch size is 8. The learning rate is $2\times 10^{-4}$, and the number of training steps is 200K.
@@ 5.2 Visualizations of Results
Following the Odyssey training process, we visualize the estimation results of SceneTracker on the LSFOdyssey test dataset, as shown in Fig. 3. We present the 12 frames of point clouds from the 40-frame test data sample at equal intervals. The predicted 3D trajectories are depicted in blue on their corresponding point clouds. From Fig. 3, we can see that our method can consistently deliver smooth, continuous, and precise estimation results in the face of the complex motions of the camera and dynamic objects in the scene.
5.3 Comparisons with the SF baseline
TABLE I: Quantitative results on the 2D metrics within the LSFOdyssey test dataset. All data come from the Odyssey training process.
@@ 5.3 Comparisons with the SF baseline
SceneTracker (Ours) 85.07 96.71 2.01
@@
TABLE II: Quantitative results on the 3D metrics within the LSFOdyssey test dataset. All data come from the Odyssey training process.
Method $\delta_{3D}^{0.10}\uparrow$ $\delta_{3D}^{0.20}\uparrow$ $\delta_{3D}^{0.40}\uparrow$ $\delta_{3D}^{0.80}\uparrow$ $\delta_{3D}^{avg}\uparrow$ $\text{Survival}_{3D}^{0.50}\uparrow$ $\text{MAE}_{3D}\downarrow$ $\text{EPE}_{3D}\downarrow$
TAP baseline 57.52 67.72 74.45 81.24 70.23 55.76 0.388 0.573
SF baseline 71.78 84.40 92.67 97.33 86.55 93.91 0.123 0.132
SceneTracker (Ours) 88.70 95.57 98.08 99.27 95.41 98.09 0.052 0.058
TABLE III: The evaluation results of SceneTracker on the LSFDriving dataset after the Odyssey training process, where only points with depth $<35m$ are considered for evaluation.
Experiment Variations vehicle background person
$\delta_{2D}^{avg}\uparrow$ $\delta_{3D}^{avg}\uparrow$ $\text{EPE}_{3D}\downarrow$ $\delta_{2D}^{avg}\uparrow$ $\delta_{3D}^{avg}\uparrow$ $\text{EPE}_{3D}\downarrow$ $\delta_{2D}^{avg}\uparrow$ $\delta_{3D}^{avg}\uparrow$ $\text{EPE}_{3D}\downarrow$
SF baseline - 77.51 39.61 1.262 81.15 55.13 0.717 40.74 23.01 1.523
SceneTracker +“Single” inference mode - 73.98 36.45 1.595 73.69 38.36 1.548 43.64 15.48 2.902
local 85.16 66.40 0.331 86.07 58.53 0.605 57.13 35.28 0.737
global 81.46 55.61 0.409 84.44 58.74 0.596 56.18 39.36 0.603
local & global 85.50 65.75 0.287 87.13 60.88 0.472 57.79 37.98 0.707
SceneTracker +“All” inference mode - 81.45 50.71 1.011 81.56 56.56 0.579 43.34 14.87 2.890
global 83.42 67.54 0.252 83.70 63.87 0.351 59.00 41.32 0.486
We follow RAFT [14] to design a straightforward scene flow baseline named the SF baseline. Since original scene flow methods can only address trajectory prediction problems between two frames, we adopt a chaining approach to handle long videos. Precisely, after calculating the scene flow between the first and second frames, we determine the $uv$ coordinates corresponding to the predicted position of the second frame and index the depth of the second frame to obtain the query point coordinates from the second frame to the third frame, and so on for subsequent frames.
@@
We also train the SF baseline on the LSFOdyssey dataset from scratch to ensure equity. We use the same training settings as the SceneTracker’s Odyssey training process. After training, we compare our SceneTracker with the SF baseline on the LSFOdyssey test dataset. The results are shown in Fig. 4, Table I and Table II.
We can see from Table I and Table II that our method significantly outperforms the SF baseline across all dataset metrics, both in 2D and 3D metrics. Especially on $\text{MAE}_{2D}$ and $\text{MAE}_{3D}$, SceneTracker achieves error reductions of 59.0% and 57.7%, respectively. In contrast to the SFE approach, our method employs identical query point features as initial templates tracked for each window, avoiding query point drift. This strategy effectively addresses occlusion and out-of-boundary effects, further substantiated in Fig. 4’s solid-box marked regions. The results fully demonstrate our method’s superior performance. The experimental findings substantiate the efficacy of our approach in tackling the LSFE challenge.
5.4 Comparisons with the TAP baseline
We observe that our estimated 3D trajectories can be projected onto the image plane to form a continuous 2D trajectory, which serves as the output of the TAP task. For this purpose, we construct a TAP baseline based on SceneTracker. Specifically, the TAP baseline first outputs the image plane $uv$ trajectory, then utilizes the $uv$ trajectory to index into the depth map of each frame, and finally transforms the indexed $uvd$ trajectory into the camera coordinate system.
We utilize SceneTracker trained through the Odyssey training process to provide 2D trajectories for the TAP baseline. After training, we compare our SceneTracker with the TAP baseline on the LSFOdyssey test dataset. The results are shown in Fig. 4 and Table II.
@@ 5.4 Comparisons with the TAP baseline
We can see from Table II that our approach significantly outperforms the TAP baseline on all 3D dataset metrics. Especially on $\text{MAE}_{3D}$ and $\text{EPE}_{3D}$, SceneTracker achieves error reductions of 86.6% and 89.9%, respectively. At the same time, on $\text{Survival}_{3D}^{0.50}$, SceneTracker achieves an 75.9% increase in accuracy. The significant performance gap stems mainly from the TAP baseline’s lack of robustness against depth noise interference and the inability to handle occlusion relationships in 3D space. It is worth noting that LSFOdyssey simulates fog and smoke, setting their depths as outliers directly, resulting in a depth map with a substantial amount of noise, posing significant challenges. The TAP baseline, due to its direct indexing from the depth map, is unable to detect and correct erroneous depth data, which is also substantiated in Fig. 4.
@@
The results fully demonstrate that the proposed method excels in spatial relationship handling and noise resistance, highlighting the unique challenges the LSFE task poses.
5.5 Results on WWDriving
This subsection mainly analyzes the performance of our method on real-world data. TABLE III shows the evaluation results of our SceneTracker under different inference modes on the LSFDriving dataset after the Odyssey training process. We evaluate the three types of points in LSFDriving separately. We test the “Single” and “All” inference modes. In the “Single” inference mode, the network tracks only one query point at a time, while in the “All” inference mode, it tracks all query points simultaneously. Since each sample in LSFDriving only requires tracking up to five points, even tracking all query points simultaneously (“All” inference mode) does not provide the network with rich cross-space context. Therefore, we make the network incorporate additional auxiliary points, such as local points within the vicinity of the target point and global points uniformly distributed across the entire image, to track them together.
It can be seen from Table III that adding auxiliary points can significantly enhance the network’s performance, irrespective of the inference mode. Simultaneously, under the condition of adding auxiliary points, the “All” inference mode exhibits superior performance across all 3D dataset metrics. Furthermore, we observe that errors in pedestrian samples are consistently higher than those in background and vehicle samples. This fully showcases the complexity and challenges posed by non-rigid motions.
Overall, our method demonstrates commendable performance in real-world scenes while training solely on synthetic datasets.
TABLE IV: Quantitative results about the template feature updates.
Experiment $\delta_{3D}^{avg}\uparrow$ $\text{Survival}_{3D}^{0.50}\uparrow$ $\text{MAE}_{3D}\downarrow$ $\text{EPE}_{3D}\downarrow$
No Update 95.27 98.04 0.055 0.061
SceneTracker 95.41 98.09 0.052 0.058
@@ 5.6 Ablation Studies
Figure 5: Dataset evaluation metrics versus the number of FIM’s updates at inference time.
TABLE V: Quantitative results about the sliding window size.
Window Size $\delta_{3D}^{avg}\uparrow$ $\text{Survival}_{3D}^{0.50}\uparrow$ $\text{MAE}_{3D}\downarrow$ $\text{EPE}_{3D}\downarrow$
8 93.32 97.77 0.070 0.075
12 94.88 98.04 0.058 0.062
16 95.41 98.09 0.052 0.058
20 85.81 97.23 0.099 0.116
24 81.47 95.41 0.144 0.174
Updates of the template feature We halt the template feature updates to assess the impact of the dynamic cost volume on the network’s performance. The performance degradation induced by the locking of template features in Table IV elucidates that LSFE necessitates a greater focus on addressing the variations in feature distribution brought about by long sequences in comparison to SFE.
Number of FIM’s updates We train the model with 4 FIM’s updates and evaluate with $n=4$. The same inference update number gives the best results as evaluated in Fig. 5.
Inference sliding window size In Table V, we inspect how the inference sliding window size affects model performance. We find that an inference window size $S=16$ achieves optimal performance across all dataset metrics.
@@ 6 Conclusion
In summary, we propose a new task, i.e., , long-term scene flow estimation, to capture the fine-grained and long-term 3D motion better. To effectively tackle this issue, we introduce a novel approach named SceneTracker. With detailed experiments, SceneTracker shows superior capabilities in handling 3D spatial occlusion and depth noise interference, highly tailored to the LSFE task’s needs. To boost the development of the LSFE field, we build the first real-world evaluation dataset, LSFDriving.